# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ I-C Main Contributions and Organization of the Paper
$\eta_{k}$, $\eta_{X}$ Depth similarity thresholds for cluster $k$ and the dataset ($\mathbb{S}$-based);
$\mathcal{I}_{k}$ Indices of points in $\mathbf{g}_{k}$;
$\mathcal{J}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ in $\mathcal{X}_{k}$ based on $\eta_{k}$;
$\mathcal{H}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ outside $\mathcal{X}_{k}$ based on $\eta_{X}$;
$\mathbf{I}_{k}$ Inner observations of cluster $k$;
$\mathbf{O}_{k}$ DBCA-outliers of cluster $k$;
II Depth-based Similarity Matrix
II-A Preliminaries
Serfling and Zuo [30] formulated several desirable properties of depth functions: affine invariance, maximality at center, monotonicity on rays, and vanishing at infinity. A deepest point $\mathbf{m}=\text{argmax}_{\mathbf{z}\in\mathbb{R}^{d}}D(\mathbf{x}\mid F)$ is usually referred to as a depth median. In applications, where the inherent distribution $F$ is usually unknown, depth is calculated with respect to the empirical distribution $F_{n}$ of $\mathbf{X}$, that is $D(\mathbf{z}\mid F_{n})$ is the depth of $\mathbf{z}$ with respect to the sample $\mathbf{X}$. Hereinafter, we will denote the latter as $D(\mathbf{z}\mid\mathbf{X})$ to emphasize the role of the sample on calculating the depth of the point $\mathbf{z}$. For a broader overview of various depth notions, see Mosler and Mozharovskyi [31].
with fixed depth median versions
where $\mathbf{z}$ denotes a point in $\mathbb{R}^{d}$, $\bar{\mathbf{X}}$ is the mean vector, $\mathbf{x}_{i}$ is an observation in $\mathbf{X}$, while $\hat{\mathbf{\Sigma}}$ is the estimated covariance matrix of dataset $\mathbf{X}$. Unlike DBSCAN, which uses fixed Euclidean distance for neighborhoods, DBCA employs (2) to define elliptical, affine-invariant neighborhoods. Recent work has also explored data depth to identify centroids for k-means initialization in diverse datasets [34, 35].
So far, the notion of depth serves an auxiliary role to the more traditional clustering methods and is not employed directly in examining the multimodality of data. Herein, two depth functions are considered for our DLCC algorithm: MD and spatial depth (SD). For SD [36], the sample version is
where $n$ is the total number of observations. Although SD is only invariant under similarity transformations, it is nevertheless a sensible choice in the current context due to its straightforward computation in high-dimensional spaces.
DLCC requires a similarity matrix to determine the neighbors of each observation. Traditional distance measures, such as Euclidean distance, are typically based on assumptions about the underlying cluster shapes (e.g., spherical or linear structures), which may not hold in more complex scenarios. In contrast, data depth offers a nonparametric, robust, and adaptive alternative. As the depths (1) and (3) provide a center-outward ordering and their value ranges from 00 to $1$, the depth values of other points can be viewed as a measure of similarity to the depth median. Thus, similarity matrices can be constructed based on data depth. For this, however, one requires an approach capable of defining neighbors of any point $\mathbf{x}_{i}$ in $\mathbf{X}$, i.e., the observations in $\mathbf{X}$ that are most similar to $\mathbf{x}_{i}$. It should be clear that the usual notion of depth cannot be applied directly here, as that can only provide similarity measurements related to the global median.
We note that this new set is comprised of all the points belonging to the original sample $\mathbf{X}$ as well as of the reflection of all the points of $\mathbf{X}$ through $\mathbf{x}_{i}$. Crucially, $\mathbf{x}_{i}$ is a depth median of $\mathbf{X}_{Ri}$ as it is a center of symmetry for this set. It is also the case that $D(\mathbf{x}_{i}|\mathbf{X}_{Ri})=1$ and that $D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ can be used to quantify the similarity of any point $\mathbf{x}_{j}$ ($j\neq i$) to the point of interest $\mathbf{x}_{i}$. Subsequently, by repeatedly computing the depth values for points in $\mathbf{X}$ with respect to $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$, a $n\times n$ non-mutual depth-based similarity matrix can be defined. We denote it as $\mathbb{S}$, where $\mathbb{S}_{ij}=D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ represents the entry in the $i$th row and $j$th column. As will be seen shortly, this can cause some computational challenges.
This approach to localize the concept of depth was introduced by Paindaveine and Van Bever [37] when they introduced local depth, which was shown to yield non-spherical depth neighborhoods in some contexts. The same authors [38] used this concept and introduced an affine-invariant k-nearest neighbor classification algorithm that adapts to the local geometry of the sample and was shown to be consistent.
II-B Mahalanobis Depth based Similarity Matrix
Previous studies [39] have employed the approach of using individual observations $\mathbf{x}_{i}$, $i=1,2,\ldots,n$, in place of the mean vector, while incorporating the globally estimated covariance matrix, for depth value computations as given in (2). In other words, the center used for calculating the Mahalanobis depth is shifted to each observation, and the depth values of the remaining observations are subsequently recorded. These studies commonly employ the minimum covariance determinant (MCD) method [40] to achieve robust results. In light of the previous discussion, the mean of $\mathbf{X}_{Ri}$ can easily be seen to be $\mathbf{x}_{i}$, which leads to
$D_{\mathrm{MD}}(\boldsymbol{z}\mid\mathbf{X}_{Ri})=D_{\mathrm{MD}}^{i}(\boldsymbol{z}),$
assuming the use of a consistent covariance matrix. This implies that making use of (2) to define a neighborhood of any point is in line with the approach of Paindaveine and Van Bever.
However, the method of Paindaveine and Van Bever necessitates estimating $n$ distinct covariance matrices, substantially increasing computational load. In contrast, the DBCA approach estimates only a single global covariance matrix, which may not adequately reflect the diversity within different clusters. To mediate between these extremes, we propose a covariance matrix selection strategy that balances computational efficiency with the ability to capture the unique characteristics of each cluster more effectively than a global approach.
In the framework of GMM, it is feasible to estimate multiple covariance matrices, ideally keeping the number of such matrices less than or equal to the number of clusters — substantially fewer than $n$ — yet more than one. The covariance matrix for cluster $k$, i.e., component $k$, can be decomposed as
${\mathbf{\Sigma}}_{k}=\lambda_{k}\mathbf{\Gamma}_{k}\mathbf{\Delta}_{k}\mathbf{\Gamma}_{k}^{\prime},$ (4)
where $\lambda_{k}=|{\mathbf{\Sigma}}_{k}|^{1/d}$, $\mathbf{\Gamma}_{k}$ denotes the matrix of eigenvectors, and $\mathbf{\Delta}_{k}$ is a diagonal matrix of eigenvalues with the condition $|\mathbf{\Delta}_{k}|=1$ [41], [42]. These elements, respectively, represent size, orientation, and shape of the cluster; see [43] for details.
In our study, we focus on the EEV model from the Gaussian parsimonious clustering models (GPCM) family, wherein different clusters estimated by the EEV model share common $\lambda_{k}$ and $\mathbf{\Delta}_{k}$ — i.e., $\lambda_{k}=\lambda$ and $\mathbf{\Delta}_{k}=\mathbf{\Delta}$ — but possess distinct $\mathbf{\Gamma}_{k}$. When model performance metrics, such as the Bayesian information criterion (BIC; [44]), are similar for results with varying numbers of components, those with fewer components are preferred. We do not consider more flexible models within the GPCM family as these could result in GMMs that overfit the data, and the size of the estimated covariance matrices may be too small for constructing the similarity matrix.
If we assume clusters are all convex, we use a test based on principal component analysis (PCA; [45]) to decide on the method for constructing the similarity matrix. With the close correlation between the Mahalanobis distance and PCA, the squared Mahalanobis distance is equal to the sum of squares of standardized principal component scores [46]. If a significant proportion of variance can be explained by the leading principal components, then the Mahalanobis distance in the space spanned by these components is a suitable metric, especially in the presence of strong inter-variable correlations. In contrast, if the variables show no clear correlations, the standard Euclidean distance is deemed a more appropriate measure.
The test procedure can be summarized as follows: Let $\mathbf{vp}=\{vp_{1},vp_{2},\ldots,vp_{H},\ldots,vp_{d}\}$, which contains the proportion of variance explained by each PC from high to low, where $H$ indicates the number of PCs that explain over $5\%$ variance, but if $H=d$, then $H$ redefined to be $d-1$. If either (5) or (6) is satisfied, the dataset passes the test:
$vp_{1}$ $>$ $0.6,$ (5)
$\sum_{h=1}^{H}vp_{h}$ $>$ $0.95$ (6)
When the dataset passes the test, we use the estimated global covariance matrix. If it fails, the EEV model is brought into play, offering the benefit of estimating clusters with identical eigenvalues. The test can be repeated using the eigenvalues derived from the EEV model. If it continues to fail, indicating weak inter-variable correlations even within clusters, and the Euclidean distance becomes sufficient for constructing the similarity matrix, setting the covariance matrix in (2) as the identity matrix. However, if it succeeds, suggesting the presence of within-cluster correlations, we employ the covariance matrices estimated by the EEV model. Specifically, consider two points, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, where $\mathbf{x}_{i}$ is assigned to cluster $1$ and $\mathbf{x}_{j}$ to cluster $2$ by the EEV model. The estimated covariance matrices, $\hat{\mathbf{\Sigma}}_{1}$ for cluster $1$ and $\hat{\mathbf{\Sigma}}_{2}$ for cluster $2$, are used to compute the $i$th and $j$th rows of the similarity matrix, respectively. While both covariance matrices are of the same size and shape, their orientations differ to capture the inter-variable correlation within each cluster.
II-C Spatial Depth based Similarity Matrix
@@
A spatial-depth-based similarity matrix can be constructed from calculating $D_{\mathrm{SD}}(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ for all $i,j=1,2,\ldots,n$. One advantage of using this depth measure is that the shape of the resulting neighborhoods is determined by the local geometry of the sample and does not rely on any parameters. This is a potentially very interesting advantage, but it does entail the computationally intensive process of repeatedly calculating depth values for the $n$ sample points with respect to $n$ distinct datasets $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$. Hence, we delve into certain properties of spatial depth in the current context (working with reflected samples) and introduce an efficient approach to calculate the similarity matrix.
@@
In order to compute the spatial depth of a point $\mathbf{x}$ from (3), we note that it is necessary to find the vectors from all sample points to $\mathbf{x}$ and their corresponding norms. In any of the combined dataset $\mathbf{X}_{Ri}$, the $n$ points from the original dataset $\mathbf{X}$ remain unchanged. Consequently, we can retain a $n\times d$ matrix, denoted as $\mathbf{E}$, to prevent redundant computations. Here, the $q^{\text{th}}$ row is the sum of the unit vectors from other observations in $\mathbf{X}$ to $\mathbf{x}_{q}$, expressed as:
@@
With these, $\mathbf{\tilde{E}}_{i}$ is expressed as: $\mathbf{\tilde{E}}_{i}=\mathbf{B}_{i}-\mathbf{R}^{\prime}_{i}\mathbf{\tilde{X}}_{i}.$ Through the above procedure, $\mathbf{\tilde{E}}_{i}$ can be derived efficiently. For instance, in a $2000\times 600$ dataset, while $\mathbf{E}$ requires approximately two seconds to compute, $\mathbf{\tilde{E}}_{i}$ only takes around $0.1$ second per loop — run on a PC with Intel(R) Core(TM) i7-11700K processor (3.6 GHz) with sufficiently large physical memory, Matlab R2023a. Having obtained $\mathbf{E}$ and $\mathbf{\tilde{E}}$, the reflection spatial depth values of any $\mathbf{x}_{q}$ in $\mathbf{X}$ based on $\mathbf{X}_{Ri}$ can be computed via
@@ III Depth-based Local Center Clustering
Neighbors and Neighborhoods: Given an integer parameter $s$, the neighbors of $\mathbf{x}_{i}\in\mathbf{X}$ are defined as the set of $s-1$ points $\mathbf{x}_{q}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}$ with the highest depth values with respect to the dataset $\mathbf{X}_{Ri}$, i.e., the points with the largest $D\left(\mathbf{x}_{q}\mid\mathbf{X}_{Ri}\right)$. The neighborhood of $\mathbf{x}_{i}$, denoted as $\mathbf{N}_{i}$, is thus formed by including $\mathbf{x}_{i}$ and its $s-1$ neighbors, constituting a subset of $\mathbf{X}$ with size $s$.
Local Depth [37] and Local Ranks: The local depth of an observation $\mathbf{x}_{i}$ is its depth value within its neighborhood $\mathbf{N}_{i}$. The local rank of an observation $\mathbf{x}_{j}$ within the neighborhood $\mathbf{N}_{i}$ is defined as:
@@
If $r_{i}(\mathbf{x}_{j})=1$, then $\mathbf{x}_{j}$ is characterized as the center of the subset $\mathbf{N}_{i}$. We define central points of $\mathbf{N}_{i}$ as a local center. These local centers can be understood as representative points for the entire dataset $\mathbf{X}$.
Figure 3: Local center example ($s=50$) for the first two PCs of the Iris dataset, where a focus point (green triangle) with its neighbors (pink points) and the corresponding local center (blue diamond) are highlighted.
@@
The DLCC method uncovers clusters through analyzing and grouping local centers. However, as $n$ subsets might yield a large number of local centers, many of them might not be useful for clustering. For example, an outlier residing between two clusters can be the deepest point in a subset containing observations from both clusters. Hence, it is crucial to filter out unnecessary local centers before grouping them. The remaining centers after filtering are denoted as filtered centers.
DLCC introduces two strategies designed to address diverse types of clustering challenges. The conditions for grouping the filtered centers under these two strategies are summarized as follows:
@@
Let ${c_{1},\ldots,c_{A}}$ denote a finite set of filtered centers and ${\mathbf{N}_{c_{1}},\ldots,\mathbf{N}_{c_{A}}}$ their corresponding neighborhoods. Define the similarity between any two neighborhoods as the proportion of overlapping observations among them. Two grouping strategies are as follows:
@@
[…]
  • Max strategy: Given a parameter $\delta$, termed as the similarity threshold, $\forall$ $c$ in group $\mathbf{g}$, $\exists c^{*}$ in the same group s.t. $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{*}})\,>\delta$, and for $\forall c^{\dagger}\notin\mathbf{g}$, $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})\,\leq\delta$.
Note that, “group” refers specifically to a set of filtered centers and should not be confused with “cluster”. Some insights into the strategy chosen for grouping local centers are important in applications. Under the min strategy, DLCC attempts to find centroids of clusters from filtered centers. The variance in the number of filtered centers across different groups allows for better flexibility than traditional center-based clustering methods, especially when dealing with clustering problems that have overlapping regions. The min strategy tends to be more suitable when the sizes of all clusters are comparable.
The max strategy assumes each cluster can be built by connecting multiple finite convex shapes created by filtered centers. Filtered centers in the same group are connected with a chain of similarities. For example, if two filtered centers $c_{y}$ and $c_{z}$ are in the same group, then there must exist a chain of filtered centers, $c_{p_{1}},c_{p_{2}},\ldots,c_{p_{h}}$, where $c_{p_{1}}=c_{y}$, $c_{p_{h}}=c_{z}$, and the similarity between any two adjacent centers’ neighborhoods in that chain is larger than $\delta$. Generally, the parameter $s$ is small under the max strategy. When using Mahalanobis depth to construct the similarity matrix with a small $s$, the max strategy is less sensitive to variations in the matrix. In such cases, a globally estimated similarity matrix often suffices, similar to DBCA. However, unlike DBCA, which builds clusters with chains of fixed-shape ellipses centered at each observation, DLCC with the max strategy considers only chains of filtered centers, with a fixed number of neighbors rather than a fixed region. Like density-based methods, the max strategy works well for disjoint clusters and can handle non-convex clusters of varying sizes.
@@
With these temporary clusters, it is expected that most observations have been allocated to a particular cluster. Nevertheless, there may still be some points that remain unclustered. In such cases, the clustering problem can be reframed as a classification problem. Intuitively, classification methods can be used to assign the remaining points to a cluster. It should be noted that the choice of classification method is very flexible here. Indeed, we could potentially consider any classification method at this stage, depending on the specific requirements of the task at hand. DLCC incorporates three built-in classification methods: maximum-depth classifier (MDC), K-nearest neighbors (KNN) based on the similarity matrix, and random forests (RF).
Lastly, self-improvement procedures of DLCC are discussed. Although the clustering problem is converted to a classification problem in the final stage, it has some special features, because the observations that are already labelled may still be updated. Two logical parameters are introduced here:
  • maxdepth: This parameter dictates whether DLCC will continue iterations until all observations in the assigned cluster attain the highest depth values.
  • ifloop: This parameter, only applicable under the min strategy, determines whether only a single filtered center should remain in each cluster.
Let the $K$ clusters constructed so far be denoted by $\mathcal{X}_{k}$, where $k=1,2,\ldots,K$. For any observation $\mathbf{x}_{j}$, $D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$ corresponds to the depth of this observation within cluster $\mathcal{X}_{k}$. If the depth value of $\mathbf{x}_{j}$ in the cluster it is currently assigned to does not match $\max_{k}D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$, and if the “maxdepth” parameter is set to TRUE, $\mathbf{x}_{j}$ will be relocated. This procedure loops until all observations are assigned to the cluster where they obtain the largest depth value. Regarding the min strategy, it can be adapted to resemble a centroid-based clustering method. If the “ifloop” parameter is TRUE, the clustering algorithm updates clusters using only the deepest filtered center within each cluster. However, other filtered centers are not disregarded. They may become the deepest center for a cluster in future iterations, as clustering results are reassessed after each update. This iterative process continues, regenerating clusters and potentially re-evaluating the roles of filtered centers, until the clustering outcome stabilizes.

### v2

@@ I-D Main Contributions and Organization of the Paper
II Spatial Depth based Similarity Matrix
@@
A spatial-depth-based similarity matrix can be constructed from calculating $D_{\mathrm{SD}}(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ for all $i,j=1,2,\ldots,n$. One advantage of using this depth measure is that the shape of the resulting neighborhoods is determined by the local geometry of the sample and does not rely on any parameters. This is a potentially very interesting advantage, but it does entail the computationally intensive process of repeatedly calculating depth values for the $n$ sample points w.r.t. $n$ distinct datasets $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$. Hence, we delve into certain properties of spatial depth in the current context (working with reflected samples) and introduce an efficient approach to calculate the similarity matrix.
@@
In order to compute the spatial depth of a point $\mathbf{x}$ from (3), we note that it is necessary to find the vectors from all sample points to $\mathbf{x}$ and their corresponding norms. In any of the combined datasets $\mathbf{X}_{Ri}$, the $n$ points from the original dataset $\mathbf{X}$ remain unchanged. Consequently, we can retain a $n\times d$ matrix, denoted as $\mathbf{E}$, to prevent redundant computations. Here, the $q$th row is the sum of the unit vectors from other observations in $\mathbf{X}$ to $\mathbf{x}_{q}$, expressed as:
@@
With these, $\mathbf{\tilde{E}}_{i}$ is expressed as: $\mathbf{\tilde{E}}_{i}=\mathbf{B}_{i}-\mathbf{R}^{\prime}_{i}\mathbf{\tilde{X}}_{i}.$ Through the above procedure, $\mathbf{\tilde{E}}_{i}$ can be computed efficiently. For example, on a $2000\times 600$ dataset, computing $\mathbf{E}$ takes approximately $2$ seconds, whereas each $\mathbf{\tilde{E}}_{i}$ requires about $0.1$ seconds per loop. All experiments in this paper were conducted on a PC equipped with an Intel(R) Core(TM) i7-11700K CPU (3.6 GHz) and sufficient physical memory, using Matlab R2023a. Having obtained $\mathbf{E}$ and $\mathbf{\tilde{E}}$, the reflection spatial depth values of any $\mathbf{x}_{q}$ in $\mathbf{X}$ based on $\mathbf{X}_{Ri}$ can be computed via
@@
From a theoretical perspective, computing depth-based similarities requires $O(n^{3})$ operations, since evaluating spatial depth for $n$ points incurs $O(n^{2})$ cost and is repeated over $n$ reflected datasets. While the proposed implementation does not reduce this complexity, it substantially accelerates computation in practice through efficient reuse of the distance relationship and matrix-level computations.
Figure 2 illustrates the empirical runtime comparison on synthetic datasets with fixed dimension $d=500$ and sample size ranging from $60$ to $3600$, reported on a logarithmic scale. For a fair comparison, we implemented the baseline in Matlab following the same algorithmic logic as the local depth in the DepthProc R package [42]. As shown, the proposed approach consistently achieves significant runtime reductions.
Figure 2: Running time comparison (log scale) between the proposed method and the traditional construction.
@@ III Depth-based Local Center Clustering
Neighbors and Neighborhoods: Given an integer parameter $s$, the neighbors of $\mathbf{x}_{i}\in\mathbf{X}$ are defined as the set of s1s-1 points $\mathbf{x}_{q}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}$ with the highest depth values w.r.t. the dataset $\mathbf{X}_{Ri}$, i.e., the points with the largest $D\left(\mathbf{x}_{q}\mid\mathbf{X}_{Ri}\right)$. The neighborhood of $\mathbf{x}_{i}$, denoted as $\mathbf{N}_{i}$, is thus formed by including $\mathbf{x}_{i}$ and its s1s-1 neighbors, constituting a subset of $\mathbf{X}$ with size $s$.
Local Ranks: The local rank of an observation $\mathbf{x}_{j}$ within the neighborhood $\mathbf{N}_{i}$ is defined as:
@@
If $r_{i}(\mathbf{x}_{j})=1$, then $\mathbf{x}_{j}$ is characterized as the center of the subset $\mathbf{N}_{i}$. We define the central point of $\mathbf{N}_{i}$ as a local center. These local centers can be understood as representative points for the entire dataset $\mathbf{X}$.
Figure 3: Local center example ($s=50$) in the two–PC visualization of the Iris dataset (see Section V for dataset description). A focus point (green triangle), its neighbors (pink points), and the resulting local center (blue diamond) are highlighted.
@@
The DLCC method uncovers clusters through analyzing and grouping local centers. However, as $n$ subsets might yield a large number of local centers, many of them might not be useful for clustering. For example, an outlier residing between two clusters can be the deepest point in a subset containing observations from both clusters. Hence, it is crucial to filter out unnecessary local centers before grouping them. The local centers that remain after going through this filtering process are denoted as filtered centers. The filtering of local centers will be discussed in details in Section III-A.
DLCC introduces two strategies for grouping filtered local centers, designed to address different types of clustering challenges. The conditions for grouping the filtered centers under these two strategies are summarized as follows.
@@
Let $c_{1},\ldots,c_{A}$} denote a finite set of filtered centers and $\mathbf{N}_{c_{1}},\ldots,\mathbf{N}_{c_{A}}$} their corresponding neighborhoods. Define the similarity between any two neighborhoods as the proportion of overlapping observations among them. For any two neighborhoods $\mathbf{N}_{c}$ and $\mathbf{N}_{c^{\prime}}$, define their similarity as
$\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\prime}})=\frac{|\mathbf{N}_{c}\cap\mathbf{N}_{c^{\prime}}|}{s},$ (6)
where $|\cdot|$ denotes set cardinality and $s$ is the neighborhood size. A grouping is a partition $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$ of the centers $\{c_{1},\ldots,c_{A}\}$. Two grouping strategies are defined as follows:
@@
    $\min_{c^{*}\in\mathbf{g}_{i}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{*}})$ $\!\!\!\!>\!\!\!\!$ $\max_{j\neq i}\min_{c^{\dagger}\in\mathbf{g}_{j}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})$ (7)
    $\max_{c^{*}\in\mathbf{g}_{i}\setminus\{c\}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{*}})$ $\!\!\!\!>\!\!\!\!$ $\max_{c^{\dagger}\notin\mathbf{g}_{i}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}}).$ (8)
  • Max strategy: Given a parameter $\delta$, termed as the similarity threshold, $\forall$ $c$ in group $\mathbf{g}$, $\exists c^{*}$ in the same group s.t. $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{*}})\,>Q$, and for $\forall c^{\dagger}\notin\mathbf{g}$,
    $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})\,\leq\delta.$
Note that, “group” refers specifically to a set of filtered centers and should not be confused with “cluster”. Some insights into the strategy chosen for grouping local centers are described in Section IV-A.
@@
With these temporary clusters, it is expected that most observations have been allocated to a particular cluster. Nevertheless, there may still be some points that remain unclustered. In such cases, the clustering problem can be reframed as a classification problem. Intuitively, classification methods can be used to assign the remaining points to a cluster. It should be noted that the choice of classification method is very flexible here. Indeed, we could potentially consider any classification method at this stage, depending on the specific requirements of the task at hand. DLCC incorporates two built-in classification methods: $k$-nearest neighbors (kkNN) based on the similarity matrix $\mathbb{S}$, and random forests (RF).